OpenStudy (anonymous):

someone please help if x+1/x=2cosa then prove x^n+(1/x)^n=2cosna please prove this using de moivre's theorem(complex number)

OpenStudy (anonymous):

First one can show that |x| =1 \[ \alpha = 2 \cos(a),\quad -2 \le \alpha \le 2 \\ x +\frac 1 x = \alpha\\ x^2 - \alpha \ x+1 =0\\ x= \frac {\alpha \pm i\ \sqrt{4 -\alpha^2}}{2}\\ |x|^2 = \frac {\alpha^2}{4} + \frac { 4 -\alpha^2}4=1\\ |x| =1 \]

OpenStudy (anonymous):

So \[ x= \cos(t) + i \sin(t)\\ \frac 1 x=\cos(t) - i \sin(t)\\\\ x + \frac 1 x = 2 \cos(a)\\ 2 \cos(t) = 2 \cos(a)\\ t= a \]

OpenStudy (anonymous):

\[ x^n = \cos( nt) + i \sin(n t)\\ \frac 1 {x^n}= \cos( nt) - i \sin(n t)\\ x^n+ \frac 1 {x^n}= 2 \cos( n t)= 2 \cos(na) \]